Please, solve this system using elimination.
3x + 5y = -18
4x - 10y = -21

Answer by mananth(16946)

3 x + 5 y = -18 .............1
4 x -10 y = -21 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
6 x 10 y = -36
4 x -10 y = -21
Add the two equations
10 x = -57
/ 10.00
x = -5.70
plug value of x in (1)
3.00 x + 5.00 y = -18.00
-17.10 + 5.00 y = -18.00
5.00 y = -18.00 + 17.10
5.00 y = -0.90
y = -0.18